20240902
Time
1955-
Attendance
Tara, Percy, Ian, Roy, Dr. Meyer
Spaces Occupied
401, Facility, Roof
Goals
- Test PICSI2
Equipment Used
Scope | Camera |
---|---|
Main | ASI 432 |
Finder | ASI 178 |
Frame Table
Time Observed | Source Name | Filter | Int Time [s] | Gain | Airmass | Mag (filter) | RA [h] | DEC [deg] | Source Type | Scope | Purpose | Notes |
---|---|---|---|---|---|---|---|---|---|---|---|---|
202409022250 | WR133 | UBVRI | e.g., 3 | 150 | 1.0 | 1.25(V) | 1.209 | 45 | Star | Main | Captures | n/a |
202409022212 | LAWD 73 | UBVRI | e.g., 3 | 150 | 1.0 | 1.25(V) | 1.209 | 45 | Star | Main | Captures | n/a |
202409022342 | MKN501 | UBVRI | e.g., 3 | 150 | 1.0 | 1.25(V) | 1.209 | 45 | Star | Main | Captures | n/a |
Log
- Roy and Tara removed ZWO Filter Wheel and 432 from on-axis
- Roy and Tara set up
- 2000 - Ian arrived and attached PICSI2 to the telescope
- 2105 - Dr. Meyer arrived
- 2105-2230 - Troubleshooting (connectivity issues with the small computer connected to PICSI2)
- Ended up running an ethernet cable all the way from the telescope down into the control room
- 2235 - Pointing using Vega
- 2255 - Observing WR133
- 2315 - Observing LAWD 73
- 2330 - Observed MKN 501
ToDo
Notes
- Need the ethernet box opened for the NUC
Weather
2% avg. sky cover